# GS-ProCams: Gaussian Splatting-based Projector-Camera Systems

### v1

@@ Abstract
We present GS-ProCams, the first Gaussian Splatting-based framework for projector-camera systems (ProCams). GS-ProCams significantly enhances the efficiency of projection mapping (PM) that requires establishing geometric and radiometric mappings between the projector and the camera. Previous CNN-based ProCams are constrained to a specific viewpoint, limiting their applicability to novel perspectives. In contrast, NeRF-based ProCams support view-agnostic projection mapping, however, they require an additional colocated light source and demand significant computational and memory resources. To address this issue, we propose GS-ProCams that employs 2D Gaussian for scene representations, and enables efficient view-agnostic ProCams applications. In particular, we explicitly model the complex geometric and photometric mappings of ProCams using projector responses, the target surface’s geometry and materials represented by Gaussians, and global illumination component. Then, we employ differentiable physically-based rendering to jointly estimate them from captured multi-view projections. Compared to state-of-the-art NeRF-based methods, our GS-ProCams eliminates the need for additional devices, achieving superior ProCams simulation quality. It is also 600 times faster and uses only 1/10 of the GPU memory.
@@
Figure 1: GS-ProCams. We represent the scene using 2D Gaussian, allowing us to efficiently model the geometric and photometric mappings of ProCams. By projecting various patterns onto the target surface and capturing images from multiple viewpoints using a camera, we achieve high-fidelity ProCams simulation while estimating the target surface’s geometry and materials. After training on a scene, GS-ProCams can achieve efficient view-agnostic (a) projector compensation, (b) projection-based diminished reality [29, 45] (with inpainting models [65]), and (c) projection mapping [28] (with diffusion models [53, 21]).
1 Introduction
@@
Projector-camera systems (ProCams) are integrated systems that combine a projector and a camera, and have emerged as indispensable tools in computer vision [60, 43, 5, 76, 3, 44, 61], spatial augmented reality (SAR)/projection mapping (PM) [66, 57, 48, 30, 13, 56, 62, 67, 22, 12, 14, 15], providing exceptional experiences in immersive interaction, artistic creation and exhibition, as well as industrial design and manufacturing, among other domains.
@@
Geometric and photometric mappings between the projector and the camera, which are foundational to ProCams, serve as cornerstones for numerous applications. For example, structured light utilizes the geometric and photometric constraints of ProCams to estimate the object geometry and materials [55, 54, 18, 39]. Furthermore, by ProCams geometric correction [6, 69, 52, 78, 75] and photometric compensation [74, 37, 19, 2, 23, 47], full projector compensation can be achieved to rectify projections on non-planar and textured surfaces, ensuring optimal visual effects.
Although recent advancements in Convolutional Neural Network-based schemes [33, 24, 32, 26, 72, 73] have greatly improved the convenience and quality of projection mapping, these methods are usually limited to a specific viewpoint, making them incapable of direct application to novel viewpoints. Neural fields-based method [14] utilize multilayer perceptrons (MLPs) [7, 64, 83] for projection mapping. Despite their capability for novel viewpoint ProCams applications, this method requires an additional co-located light and a dark room (i.e., assumes no ambient light). Moreover, it suffers from low computational and memory efficiency.
In this paper, we propose GS-ProCams for efficient view-agnostic projection mapping. In particular, we first utilize 2D Gaussians [27] as ProCams scene representation, since the original 2D Gaussians do not explicitly capture Bidirectional Reflectance Distribution Function (BRDF), we augment each Gaussian point in space with BRDF attributes. Then, we capture images of the target surface within the scene from multiple camera viewpoints while using a fixed projector to project various patterns onto it. Afterward, we employ structure-from-motion (SfM) [63, 58] to determine the poses and intrinsic parameters of the projector and the cameras. Subsequently, we jointly reconstruct the target surface and the projector’s response by differentiable physically-based rendering across each viewpoint. Finally, we apply this trained GS-ProCams model to various view-agnostic ProCams applications, such as projection mapping and projector compensation, as shown in Fig. 1. Comprehensive experiments on ProCams simulation and projector compensation show clear advantages of our GS-ProCams over previous methods.
@@
  • Compared to NeRF-based ProCams, GS-ProCams eliminates the need for additional devices, can operate under ambient room light, and is 600 times faster while using only 1/10 of the GPU memory.
@@
  • We introduce a view-agnostic ProCams benchmark encompassing various viewpoints, textured surfaces, and ambient lighting conditions for training and evaluating ProCams models. This benchmark is designed to foster future research in the field.
@@ 2.1 ProCams simulation
ProCams simulation aims to model the real-world photometric and geometric mappings of projector-camera systems (ProCams) within a computational space. Once these mappings are established, the simulation synthesizes camera images corresponding to various projections. ProCams simulation facilitates image-based relighting [74, 24, 14], shape reconstruction [60, 43, 24], and projector compensation [32, 26]. Early methods describe the pixel mapping from projectors to cameras using the Light Transport Matrix (LTM) [59, 49, 50, 74]. However, the reconstruction quality of LTM, which is based on the linear relationship between pixels, generally depends on dense sampling and specialized computational designs of large matrices [10, 11, 59, 74, 49, 70]. By employing neural rendering, DeProCams [24] generates photorealistic simulations of ProCams with fewer image samples by incorporating inherent epipolar geometric and simplified photometric constraints under a fixed viewpoint. However, a notable limitation of this approach is its reliance on manual calibration of ProCams using a checkerboard pattern [25], rather than leveraging data priors [58]. Although Park et al. [51] examine the geometric and photometric mappings in ProCams using a differentiable rendering scheme, it remains limited to specific view information and relies on an RGB-D camera. Additionally, the method depends heavily on pixel-level bias corrections. Erel et al. [14] integrate projectors into Neural Reflectance Fields [7, 64, 83] to enable view-agnostic projection mapping within 3D scenes. Despite this progress [14], the influence of other light sources within the scene is neglected, and a light source co-located with the camera is required, thereby causing limited applicability. We summarize the characteristics of technologies for ProCams simulation to provide an intuitive comparison with our approach in Tab. 1.
Table 1: Comparison of representative schemes for ProCams and proposed GS-ProCams. GS-ProCams is a view-agnostic method, which enables multi-view applications. It also simultaneously addresses geometric and photometric mappings under ambient light, using only an RGB camera and a projector.
Method View-agnostic W/o manual W/o extra Ambient
@@
Depth-prior DR [51]
NeRF-like DR [14]
@@ 2.2 Projector compensation
Projector compensation aims to adjust projections for non-planar and textured surfaces to achieve the best visual display effect by optimizing the projector’s input patterns. Typically, projector compensation can be conducted separately through geometric correction [6, 69, 52, 78, 75] and radiometric compensation [19, 37, 2, 23, 47]. Upon completing the ProCams simulation, the compensated projection patterns can be obtained by multiplying the desired images with the inverse LTM [74]. Differentiable ProCams [24, 51, 14] generate projection patterns by making the synthesized scene closely approximate the desired images. On the other hand, comprehensive CNN-based architectures enable end-to-end projector compensation by ingeniously integrating geometric correction estimation represented by the warping grid [26, 72] or optical flow [73, 33, 40] with another component for photometric compensation. However, these methods are typically constrained to a specific viewpoint and smooth, uniform surfaces. Erel et al. [14] achieve projector compensation from novel viewpoints by integrating a projector into Neural Reflectance Fields [7, 64, 83]. Despite that, this method, which relies on ray-casting and MLPs, poses challenges to computational and memory resources. Additionally, this method assumes no ambient light in the scene.
@@ 2.3 Gaussian splatting and inverse rendering
3D Gaussian Splatting (3DGS) [35] has quickly sparked widespread interest due to its real-time differentiable rendering capability. While achieving rendering quality comparable to the state-of-the-art methods in novel view synthesis [42, 4, 77, 46, 16], it significantly enhances the efficiency of training and rendering. Furthermore, inverse rendering schemes extend 3DGS for materials and ambient lighting estimations [31, 17, 41].
While 3DGS [35] focuses primarily on efficient image rendering capabilities, reconstructing precise geometric shapes directly from Gaussian points poses significant challenges [27, 9, 80, 20, 81, 68]. In contrast, 2D Gaussian Splatting (2DGS) [27] uses 2D Gaussian primitives to ensure geometric consistency across multiple viewpoints. We enhance 2DGS with BRDF parameterization to capture the interactions between projector lighting and surfaces. We first use 2DGS to generate depth and materials maps in the camera frame and then perform shading calculations. The entire process is fully explicit and differentiable.
Figure 2: GS-ProCams setup. The intersection point $\mathbf{x}_{s}$ of the ray from the camera and the target surface is determined by 2D Gaussians. The direct light from the projector illuminates this point.
Figure 3: Pipeline. GS-ProCams uses 2D Gaussians to represent the scene, then obtains the geometry and materials of the target surface, as well as the global illumination component, through differentiable splatting techniques [27]. By using various projection patterns and capturing images from different viewpoints, GS-ProCams jointly optimizes the explicit projector responses and attributes of the 2D Gaussian points listed on the left through physically-based differentiable rendering.
@@ 3.1 Preliminary
The orientation of a 2D Gaussian can be construed as a rotation matrix $\mathbf{R}=[\mathbf{t}_{u},\mathbf{t}_{v},\mathbf{t}_{w}]$ and be represented in the form of a unit quaternion $\mathbf{q}$ during optimization. In summary, each 2D Gaussian is parameterized by several learnable parameters as $\{\mathbf{p},\mathbf{q},\mathbf{s},o,\mathbf{c}\}$. Furthermore, the designed density control mechanism enables adaptive regulation of the total number of points and the density within a unit volume.
@@ 3.2 ProCams modeling
The core of our proposed GS-ProCams is establishing the geometric and photometric mappings of projector-camera systems (ProCams) using Gaussian splatting, thus achieving high-fidelity ProCams simulation. Consider a point $\mathbf{x}_{s}$ on the target surface within a scene, which is captured by the camera as $\mathbf{x}_{c}$ in the camera frame, as illustrated in Fig. 2. In the absence of occlusion, this point is illuminated by $\mathbf{x}_{p}$ in the projector frame. Furthermore, $\mathbf{x}_{s}$ is affected by ambient and complex indirect lighting, i.e., global illumination. Taking these factors into consideration, We utilize the classical rendering equation [34] to describe the relationship between outgoing and incoming light at point $\mathbf{x}_{s}$ as follows:
[…]
where $\Omega$ represents the upper hemisphere centered at $\mathbf{x_{s}}$, with the normal vector $\mathbf{n}_{s}$ and $f$ denotes the BRDF. We employ 2D Gaussians described in 2D Gaussian Splatting (2DGS) [27] as foundational primitives, extending their original attributes by BRDF parameterization to capture the interactions between the light and the target surface. Building upon this, We capture a series of images of the surface $\{\mathbf{I}_{c}^{m}\}_{m=1}^{M}$ from multiple viewpoints, which is illuminated by known projector input patterns $\{\mathbf{I}_{p}^{n}\}_{n=1}^{N}$. Then, we simultaneously estimate the surface’s unknown geometric and material properties utilizing physically-based rendering (PBR). We illustrate the pipeline of our proposed GS-ProCams in Fig. 3.
@@ 3.3 Attributes from Gaussians
As shown in Fig. 3, we first derive attributes of the target surface from 2D Gaussians. The attributes can generally be divided into three categories: (1) the depth and inherent normals from geometry attributes of Gaussians; (2) the BRDF attributes for material reflectance; and (3) the global illumination component, approximated by RGB residuals. Notably, we utilize normals derived from depth gradients in the shading computations, while use inherent normals to regularize geometric consistency.
Depth. The intersections between camera rays and the surface are determined by their depths in the camera frame. We use the same approach as 2DGS [27] and render the mean depth $D$ by Eq. 3 to represent the ray-surface intersection as follows:
@@
where $\mathbf{x}$ is the homogeneous ray corresponding to the camera pixel $\mathbf{x}_{c}=(x,y)$, and $z_{i}$ denotes the depth of the intersection between $\mathbf{x}$ and the $i$-th Gaussian planes.
@@
BRDF attributes. To enable the light-surface interactions, we augment each original 2D Gaussian point [27] with additional learnable attributes: albedo $\mathbf{b}\in[0,1]^{3}$ and roughness $r\in[0,1]$. Then we splat these attributes onto the camera frame by utilizing the efficient differentiable rasterization method described by Eq. 3:

### v2

@@ Abstract
We present GS-ProCams, the first Gaussian Splatting-based framework for projector-camera systems (ProCams). GS-ProCams is not only view-agnostic but also significantly enhances the efficiency of projection mapping (PM) that requires establishing geometric and radiometric mappings between the projector and the camera. Previous CNN-based ProCams are constrained to a specific viewpoint, limiting their applicability to novel perspectives. In contrast, NeRF-based ProCams support view-agnostic projection mapping, however, they require an additional co-located light source and demand significant computational and memory resources. To address this issue, we propose GS-ProCams that employs 2D Gaussian for scene representations, and enables efficient view-agnostic ProCams applications. In particular, we explicitly model the complex geometric and photometric mappings of ProCams using projector responses, the projection surface’s geometry and materials represented by Gaussians, and the global illumination component. Then, we employ differentiable physically-based rendering to jointly estimate them from captured multi-view projections. Compared to state-of-the-art NeRF-based methods, our GS-ProCams eliminates the need for additional devices, achieving superior ProCams simulation quality. It also uses only 1/10 of the GPU memory for training and is 900 times faster in inference speed. Please refer to our project page for the code and dataset: https://realqingyue.github.io/GS-ProCams/.
keywords:
Projection mapping, projector-camera systems, gaussian splatting.
[…]
@@
GS-ProCams. We represent the scene using 2D Gaussian, allowing us to efficiently model the geometric and photometric mappings of ProCams. (A) By projecting various patterns onto the projection surface and capturing images from multiple viewpoints using a camera, (B) we achieve high-fidelity ProCams simulation while estimating the projection surface’s geometry and materials. (C) After training on a scene, GS-ProCams can achieve efficient view-agnostic projector compensation, projection-based diminished reality [28, 46] (with inpainting models [65]), and projection mapping [27] (with diffusion models [55, 20]).
@@
Introduction Projector-camera systems (ProCams) are integrated setups that combine one or more projectors and cameras, and have emerged as indispensable tools in computer vision [62, 43, 45, 63], and in spatial augmented reality (SAR)/projection mapping (PM) [27, 66, 1]. ProCams have enabled exceptional experiences in immersive interaction [12, 21, 41, 58, 14], artistic creation and exhibition [59, 52, 79, 11], as well as industrial design and manufacturing [67], among other domains. Alongside these developments, recent advances in image generation [55, 65, 82] have introduced powerful capabilities for projection mapping. These techniques enable ProCams to interactively and convincingly alter the perceived appearance of real-world objects, thereby significantly expanding the expressive and functional scope of ProCams-based applications.
@@
At the core of ProCams-based applications lie geometric and photometric mappings between the projectors and cameras. Structured light utilizes the geometric and photometric constraints of ProCams to estimate the object geometry and materials [57, 56, 17, 37]. Furthermore, through geometric registration [5, 69, 54, 77, 75] and photometric compensation [74, 35, 18, 3, 22, 48], ProCams can achieve full projector compensation to rectify projections on non-planar and textured surfaces, ensuring high-fidelity visual results.
Despite significant efforts to model projector-surface interactions under the assumption of a static spatial relationship between the projector and the projection surface, most existing methods [31, 23, 30, 24, 72, 73, 38] establish geometric and photometric correspondences that are tightly coupled to a specific viewpoint. As a result, these learned or computed relationships fail to generalize when the viewpoint changes. The neural field-based framework [13] utilizes multilayer perceptrons (MLPs) [6, 64, 84] to construct an integrated neural representation for view-agnostic ProCams simulation. While it disentangles objects’ geometry and materials through a multi-stage training strategy and supports view-agnostic projector compensation, this method requires an additional co-located light and a dark room (i.e., assumes no environment light). Moreover, it suffers from low computational and memory efficiency, which limits its applicability in real-world scenarios.
We propose GS-ProCams, an efficient and fully differentiable framework for view-agnostic projection mapping, enabling practical applications with non-static observers, as illustrated in GS-ProCams: Gaussian Splatting-Based Projector-Camera Systems. A key challenge in this domain is modeling the complex interaction between the projector and real-world surfaces. To address this, we utilize 2D Gaussians [26] as the ProCams scene representation. Since the original 2D Gaussians do not explicitly capture the Bidirectional Reflectance Distribution Function (BRDF), we augment each Gaussian point with BRDF attributes. To support dynamic, high-resolution projection patterns, we first derive surface attributes from Gaussians and then feed them into a unified lighting computation, where a compact convolution kernel is jointly optimized to capture the projector’s Point Spread Function (PSF). We further leverage the anisotropic expressiveness of 2D Gaussians to approximate residual global illumination as a view-dependent residual color, effectively capturing environment illumination and enhancing applicability. Rather than merely combining existing techniques, our framework is designed with a unified representation and optimization objective that couples the projector responses, scene geometry, material attributes, and residual illumination into a single differentiable system. Experiments on both ProCams simulation and projector compensation show clear advantages of our GS-ProCams over previous methods.
@@
  • Compared to NeRF-based ProCams, GS-ProCams eliminates the need for additional devices, can operate under ambient room lighting, and is 900 times faster in inference speed while using only 1/10 of the GPU memory for training.
@@
  • We introduce a view-agnostic ProCams benchmark encompassing various viewpoints, textured surfaces, and environmental lighting conditions for training and evaluating ProCams models. This benchmark is designed to foster future research in the field.
@@ 1.1 ProCams simulation
ProCams simulation aims to model the real-world photometric and geometric mappings of projector-camera systems (ProCams) within a computational space. Once these mappings are established, the simulation synthesizes camera-captured images corresponding to various projections. ProCams simulation facilitates image-based relighting [74, 23, 13, 38], shape reconstruction [62, 43, 23, 13], and projector compensation [30, 72]. Early methods describe the pixel mapping from projectors to cameras using the Light Transport Matrix (LTM) [61, 49, 50, 74]. However, the reconstruction quality of LTM, which is based on the linear relationship between pixels, generally depends on dense sampling and specialized computational designs of large matrices [9, 10, 61, 74, 49, 70].
DeProCams [23] generates photorealistic simulations of ProCams using fewer image samples by incorporating inherent epipolar geometric and simplified photometric constraints under a fixed viewpoint with neural rendering based on a convolutional neural network (CNN). However, a notable limitation of this approach stems from its view-specific design, which precludes the use of multi-view geometric constraints and necessitates manual ProCams calibration with a checkerboard pattern [25]. Although Park et al. [51] examine the geometric and photometric mappings in ProCams using a differentiable rendering scheme, it remains limited to specific view information and relies on an RGB-D camera. Additionally, the method depends heavily on pixel-level bias corrections. Li et al. [38] model physical ProCams using interpretable parameters based on path tracing to enable high-fidelity lighting effects, and benefit downstream ProCams tasks. However, it relies on geometry reconstructed from a specific viewpoint by gray-coded structured light (SL) [44], and is thus sensitive to the quality of this reconstruction, making it unsuitable for direct application to novel viewpoints.
Modeling view-agnostic ProCams remains a challenging task. It requires capturing the geometry of the projection surface and modeling view-dependent reflectance effects, as most real-world surfaces exhibit non-Lambertian properties. Moreover, environmental illumination is often non-negligible in practical indoor environments. For view-specific ProCams, the influence of environment light on the projection surface can be easily accounted for by capturing an image from the scene under environment light only at the same viewpoint. This can be achieved by projecting a plain black image, and the final appearance of the projection surface can then be approximated by linearly combining this with the radiance induced by the projected content. However, for novel viewpoints where synthesis is required, such information is not available beforehand. Erel et al. [13] integrate the projector into a neural reflectance field [6, 64, 84] to enable view-agnostic projection mapping within 3D scenes. Despite this breakthrough [13], the influence of other light sources in the environment is ignored, and the approach requires a light source co-located with the camera, which limits its applicability. In Table˜1, we detail the features and requirements of current ProCams simulation approaches to allow for a clear comparison with our GS-ProCams.
Table 1: Comparison of representative approaches for ProCams and our GS-ProCams. We use and to indicate the limitations and capabilities of different methods, respectively. Notably, view-specific methods can account for environmental illumination by capturing the scene under environment light only at the corresponding viewpoint. However, such information is not available in advance for unseen novel viewpoints required by view-agnostic ProCams. We use highlights to emphasize this distinction. GS-ProCams is a view-agnostic method that enables multi-view applications. It simultaneously handles both geometric and photometric mappings under environment light, using only an RGB camera and a projector.
Method View-agnostic W/o manual W/o extra Environment
@@
Depth-prior [51]
Path-tracing [38]
NeRF-based [13]
@@ 1.2 Projector compensation
Projector compensation adapts projected images to non-planar and textured surfaces to achieve the desired visual appearance. It typically involves geometric [5, 69, 54, 77, 75] and photometric [18, 35, 3, 22, 48] compensation, which have been studied separately in prior work. End-to-end methods enable projector compensation by integrating geometric correction estimation represented by the warping grid [24, 72] or optical flow [73, 31, 39] with CNN-based photometric compensation modules. While recent work [72] improves computational and memory efficiency for high-resolution projections, CNNs still struggle to preserve high-frequency details due to spectral bias [53], often resulting in oversmoothed outputs. Furthermore, these methods are typically constrained to a specific viewpoint and smooth, uniform surfaces.
An alternative strategy is to leverage ProCams simulation as a forward model to guide projector compensation. Traditional approaches [74] obtain the compensated projector input pattern by applying the inverse LTM. More recently, fully differentiable ProCams [23, 51, 13, 38] generate projection patterns by making the synthesized scene closely approximate the desired appearances. Erel et al. [13] achieve projector compensation from novel viewpoints by view-agnostic ProCams simulation using a neural reflectance field [6, 64, 84]. Despite that, this method, which relies on ray-casting and MLPs, poses challenges to computational and memory resources. Additionally, this method assumes no environment light in the scene.
@@ 1.3 Gaussian splatting and inverse rendering
3D Gaussian Splatting (3DGS) [33] utilizes 3D Gaussians to represent scenes, enabling multi-view photorealistic image rendering through rasterization techniques. It is renowned for its real-time differentiable rendering capabilities while maintaining rendering quality that rivals leading methods in novel view synthesis [42, 4, 76, 47, 15]. Additionally, recent advancements extend 3DGS to estimate materials and environment lighting within an inverse rendering framework [29, 16, 40].
Although 3DGS [33] is optimized for efficient image rendering, extracting precise geometrical details from Gaussian points is challenging [26, 8, 80, 19, 81, 68]. Conversely, 2D Gaussian Splatting (2DGS) [26] employs 2D Gaussian primitives to ensure geometric consistency across multiple viewpoints, but does not explicitly model materials and lighting. We extend 2D Gaussian points by incorporating BRDF parameterization to simulate the interaction between projector light and surfaces in ProCams.
Figure 1: GS-ProCams setup. The intersection point $\mathbf{x}_{s}$ of the ray from the camera and the projection surface is determined by 2D Gaussians. The direct light from the projector illuminates this point.
Figure 2: GS-ProCams training pipeline. GS-ProCams uses 2D Gaussians to represent the scene, then obtains the geometry and materials of the projection surface, as well as the global illumination component, through differentiable splatting techniques [26]. By using various projection patterns and capturing images from different viewpoints, GS-ProCams jointly optimizes the explicit projector responses and attributes of the 2D Gaussian points listed on the left through physically-based differentiable rendering.
@@ 2.1 Preliminary
The orientation of a 2D Gaussian can be constructed as a rotation matrix $\mathbf{R}=[\mathbf{t}_{u},\mathbf{t}_{v},\mathbf{t}_{w}]$ and represented as a unit quaternion $\mathbf{q}$ during optimization. In summary, each 2D Gaussian is parameterized by several learnable parameters as $\{\mathbf{p},\mathbf{q},\mathbf{s},o,\mathbf{c}\}$. Furthermore, the designed density control mechanism enables adaptive regulation of the total number of points and the density within a unit volume.
@@ 2.2 ProCams modeling
The core of our GS-ProCams is establishing the geometric and photometric mappings of projector-camera systems (ProCams) using Gaussian Splatting (GS), thus achieving view-agnostic and high-fidelity ProCams simulation. As illustrated in Fig.˜1, consider a point $\mathbf{x}_{s}$ on the projection surface within a scene, which is captured by the camera as $\mathbf{x}_{c}$ in the camera frame. In the absence of occlusion, this point is illuminated by the pixel $\mathbf{x}_{p}$ in the projector frame. Furthermore, $\mathbf{x}_{s}$ is also illuminated by environment and complex indirect lighting, i.e., global illumination. Taking these factors into consideration, we utilize the classic rendering equation [32] to describe the relationship between outgoing and incoming light at point $\mathbf{x}_{s}$ as follows:
$L_{o}(\omega_{o},\mathbf{x}_{s})=\int_{\Omega}f(\omega_{o},\omega_{i},\mathbf{x}_{s})L_{i}(\omega_{i},\mathbf{x}_{s})(\omega_{i}\cdot\mathbf{n}_{s})d\omega_{i},$ (4)
where $\Omega$ represents the upper hemisphere centered at $\mathbf{x_{s}}$, with the normal vector $\mathbf{n}_{s}$, and $f$ denotes the BRDF. We employ 2D Gaussians described in Section˜2.1 as foundational primitives, extending their original attributes by BRDF parameterization to capture the interactions between the projector’s direct light and the projection surface. As the projector’s direct illumination dominates the scene, we estimate a view-dependent residual color term to account for global illumination, as detailed in Section˜2.4, avoiding the computational cost of full global illumination modeling as in DPCS [38]. Building upon this, we capture a series of images of the surface $\{\mathbf{I}_{c}^{m}\}_{m=1}^{M}$ from multiple viewpoints, which is illuminated by known projector input patterns $\{\mathbf{I}_{p}^{n}\}_{n=1}^{N}$. Then, we simultaneously estimate the surface’s unknown geometric and material properties utilizing physically-based rendering (PBR). We illustrate the pipeline of our GS-ProCams in Fig.˜2.
@@ 2.3 Attributes from Gaussians
As shown in Fig.˜2, we first derive attributes of the projection surface from 2D Gaussians positioned in 3D space. The attributes can generally be divided into three categories: (1) the depth and inherent normals from geometry attributes of Gaussians; (2) the BRDF attributes for material reflectance; and (3) the global illumination component, approximated by view-dependent RGB residual color. Notably, we utilize normals derived from depth gradients in the shading computations, while using inherent normals to regularize geometric consistency.
Depth. The intersections between camera rays and the surface are determined by their depths in the camera frame. We use the same approach as 2DGS [26] and render the mean depth $D$ by Eq.˜3 to represent the ray-surface intersection as follows:
@@
where $z_{i}$ denotes the depth of the intersection between the homogeneous ray $\mathbf{x}$ corresponding to the camera pixel $\mathbf{x}_{c}=(x,y)$ and the $i$-th Gaussian planes.
@@
BRDF attributes. To enable the light-surface interactions for the projector, we augment each original 2D Gaussian point [26] with additional learnable attributes: albedo $\mathbf{b}\in[0,1]^{3}$ and roughness $r\in[0,1]$. Then we splat these attributes onto the camera frame by utilizing the efficient differentiable rasterization method described by Eq.˜3: